OpenStudy (anonymous):

Sketch the graph of the given function. g(x) = x3 − 3x, domain [−2, 2] So I have a small question on this problem. I am finding the x-values, so I set the equation=0 and I got x=0,x=+sqrt 3, x=-sqrt 3. I am not sure of the y coordinates to each of these values, any ideas?

OpenStudy (anonymous):

what does the function look like

OpenStudy (anonymous):

can you see this link? http://www.webassign.net/waplots/9/0/918550fe567c22adff6204a23facdd.gif

OpenStudy (anonymous):

well example if x=0 then I have y=g(x)=0^3+3(0)=0 next is x=3 then I have y=3^3+3(3)=27+9=36

OpenStudy (anonymous):

sorry thats plus so 27-9=18

OpenStudy (anonymous):

I thought you're just supposed to set the whole equation=0 to find the x-intercepts? x3 − 3x=0

OpenStudy (anonymous):

then \[x ^{3}-3x= x(x ^{2}-3) \] so x=0 or x\[x ^{2}-3=0\] so x=[\sqrt{3}\]

OpenStudy (anonymous):

right

OpenStudy (anonymous):

\[x=\pm \sqrt{3}\]

OpenStudy (anonymous):

I am just not sure what the y value would be..would I write it as sqrt3,0 or something? or is there a y value? because I tried that and it didnt work

OpenStudy (anonymous):

because out of this equation there are 3 x-values, x=0, x=sqrt 3, x=sqrt -3

OpenStudy (anonymous):

the question says to graph, is it asking for x cooridants

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

sorry I already graphed it

OpenStudy (anonymous):

there's just multiple parts to the question

OpenStudy (anonymous):

and I am just stuck on the coordinates part

OpenStudy (anonymous):

what is the other parts?

OpenStudy (anonymous):

I mean I got the x's just dont know what the y's would be

OpenStudy (anonymous):

Indicate the x- and y-intercepts.

OpenStudy (anonymous):

so there are three of them like I said earlier but Idk how I would write it; like (sqrt3,0) (whats the y?)

OpenStudy (anonymous):

on x intercepts, the y is 0. so write it as the pair (value, 0)

OpenStudy (anonymous):

I tried that it didnt work :/

OpenStudy (anonymous):

well (0,0) and (-sqrt3,0) and (sqrt3,0)

OpenStudy (anonymous):

yea I did that for some reason its saying its wrong.

OpenStudy (anonymous):

please copy the question and let me see

OpenStudy (anonymous):

nevermind I got it! I guess this program I am working with just requires a different order. thanks !